OpenStudy (anonymous):

integrate (1+ cosx)1/2

OpenStudy (anonymous):

\[\sqrt{1+ \cos x}\] do you mean this?

OpenStudy (anonymous):

why we root the equation

OpenStudy (anonymous):

does the question look like this \[\sqrt{1+ \cos x}\] or is it just that (1+ cos x) is multiplied by half?

OpenStudy (anonymous):

it is (1+ cos x) is multiplied by half

OpenStudy (anonymous):

x/2 +sinx/2

OpenStudy (anonymous):

then ishaan is correct

OpenStudy (anonymous):

but how i need to know i cant understand

OpenStudy (anonymous):

is it \[\int \frac{1+\cos(x)}{2}dx\] or \[\int (1+\cos(x))^{\frac{1}{2}}dx\] b money is on the second one

OpenStudy (anonymous):

1/2 + cos x/2 dx integrating the first expression simply gives x/2 and the second sinx/2 there is no complicated algebra...just the basic formulas to integrate are used. check the list of formulas in your book for cos x, you will find that its sin x.

OpenStudy (anonymous):

and integration of dx is simply x.

OpenStudy (anonymous):

my money is still on the second one

OpenStudy (anonymous):

especially in light of next question.

OpenStudy (anonymous):

she confirmed its the first one. she might have deleted that post.